OpenStudy (anonymous):

A styrofoam Snowman Holiday decoration is made by connecting 3 different "spheres" with toothpicks. The top sphere has a radius of 1", the middle sphere has a radius of 2", and the bottom sphere has a radius of 3". To the nearest cubic inch, what is the volume of this Snowman? 216 cubic inches 150 cubic inches 36 cubic inches 81 cubic inches

OpenStudy (anonymous):

I picked 150.

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

You're right :)

OpenStudy (anonymous):

volume = 4/3 pi r^3 v1 + v2 + v3 = 4/3pi (1 + 8 + 27) = 4/3 pi (36) = 150 cubic in

OpenStudy (anonymous):

Thanks!